In a class of 31 students , 22 study physics , 20 study chemistry & 5 study neither, how many students in the class study both
Answer by stanbon(75887)

In a class of 31 students , 22 study physics , 20 study chemistry & 5 study neither, how many students in the class study both
----------------
Draw a Venn Diagram with 2 intersecting circles labeled physics and chemistry.
----
Let "x" be the number who study both::
Put "x" in the intersection of the 2 circles.
Then 22-x = # who study physics but not chem.
And 20-x = # who study chem and not physics.
Note: 5 are in the rectangle but not in either circle.
---
Equation:
# in circles:: 22-x + 20-x + x = 31-5
42-x = 36
x = 6 (# in the class who study both)
-------------------------
Cheers,
Stan H.
--------------------